Wednesday, March 31, 2010
Tuesday, March 30, 2010
Around 20:00 UT Monday the AIA CCD heaters were turned off allowing the CCD to temperatures drop rapidly from about +40C to -70C and then slowly settle towards their final temperatures. With that drop in temperature, we saw the camera background decrease markedly, and the image quality in all channels improve dramatically.
The engineering images are beautiful, even though taken in an approximate focus position. Tomorrow, we plan to make a series of focus scans to determine an initial optimal focus, and work will start on calibrating the detector amplifier gains, instrument stabilization system response, etc., which will continue to improve image quality.
Monday, March 29, 2010
Saturday, March 27, 2010
The AIA CCDs are still warm. On Monday the decontamination heaters will be turned off, and the sequencer will be started so the CCD cool-down can be observed.
All nine SDO instrument doors are now open.
Friday, March 26, 2010
Thursday, March 25, 2010
Wednesday, March 24, 2010
At 11:00 am EST this morning HMI began opening their door. This door protected the front window and optics from damage but can now be opened to let the Sun shine in. Sunlight was used to illuminate the front window and study the optical performance of the telescope. A sunspot could be seen in these engineering images.
Tuesday, March 23, 2010
The instrument teams of SDO began their work to start taking data. HMI allowed its CCDs to cool over the weekend and EVE started the cooldown cycle on Tuesday. AIA plans to watch the Sun while their CCDs cool to operational temperatures later this week. The Ka-band transmitter was turned on and high-rate (150 Mbps) data is flowing to the SOCs.
Saturday, March 20, 2010
The EVE instrument completed its component check outs.
HMI prepared to turn off the decontamination heaters on Saturday to allow the CCDs to get cold.
AIA worked with their guide telescopes and image stabilization system. AIA'S GTs tell us where we are pointing with great precision and they are important to test.
SDO continues to operate well. Another longer eclipse will happen tonight and an antenna handover will be done.
Friday, March 19, 2010
doing very well !!!
We turned on ESP, MEGS-A CCD, and MEGS-B CCD for this first time this
morning. Their dark data look fine. Every component looks healthy.
Congratulations and thanks to the EVE team and GSFC SDO project team !!!!"
Thursday, March 18, 2010
Tuesday, March 16, 2010
Sunday, March 14, 2010
Friday, March 12, 2010
Wednesday, March 10, 2010
Tuesday, March 9, 2010
Now that we are close to our final orbit the main engine plumbing will be vented to exhaust fuel and oxidizer from the lines. The first Trim Motor Firing (TMF #1) is planned for Thursday evening.
Saturday, March 6, 2010
A successful Apogee Motor Firing #7 continued the string of smooth Main Engine burns since we removed the structural filter and lengthened the settling burn. AMF-7 raised our perigee altitude to approximately 25000 km, giving SDO an orbital period of about 19.2 hours. This sets SDO for one last trip "around the world" before Monday's AMF-8 actually moves SDO into a nearly geosynchronous orbit. After AMF-8, there will be three additional Trim Motor Firings (TMFs) to fix our operational orbit over White Sands.